# TrajAgent: An Agent Framework for Unified Trajectory Modelling

### v1

@@ 2.3. AutOpt
Due to the geospatial heterogeneity and the diversity of trajectory data, the trajectory models usually cannot be directly transferred between data and regions. In other words, for various data in different region, the model needs to be trained from scratch. Thus, the sufficient optimization of various models becomes emergent. In TrajAgent, we design AutOpt module to complete this task. As shown in the right part of Figure 1, AutOpt provides several optimization mechanisms: data augmentation, parameter optimization, joint optimization, prompt optimization and optimization via tools. Following Reflexion (Shinn et al., 2024), we design optim agent for each kind of optimization mechanism. The standard optim agent utilizes the history operation and related results as the feedback to update its action in the next step. Specifically, it works as two stages, including ”think” and ”action”. To support the ”think then action”, it builds a long-term memory for recording all experimental data and a short-term memory for historical actions. In the ”think” stage, optim agent analyzes the long-term memory and meta information of experiment and generating the guidance of action in the short-term memory. In the ”action” stage, optim agent analyze the results in short-term memory to generate the action. Different optimization mechanism utilize the same optim agent with different action space and optimization tips.
2.3.1. Data Augmentation
@@
Based on the standard optim agent, we introduce the specific action space for data augmentation. Following practice from existing works (Zhou et al., 2024; Dang et al., 2023; Yin et al., 2024), we define a fixed set with ten operators for data augmentation, e.g., insert, replace, split and so on. During the optimization, the operator set with simple description is provided to the optim agent, it can select any combination of operators with their simple parameters as the action. After action, optim agent obtain the feedback information, e.g., performance metrics, from UniEnv to continue update its memory and action.
2.3.2. Parameter Optimization
@@
The action space of parameter optimization is defined based on the parameters of model itself. We define a parameter configuration file for each model, the optim agent reads the configuration file and generates code as the action to update the parameters in it. To better understand the meaning of each parameter, we add comments for each parameter in the file. While the action space is large, this kind of action space is flexible to adapt with any models.
2.3.3. Joint Optimization
@@
Furthermore, we introduce the joint optimization mechanisms in AutOpt to improve the final performance. Due to the different working paradigms, the direct combination of two kinds of optimizations is unsuccessful. We designate the optimization order to prioritize data augmentation first, followed by parameter optimization. This means that once the performance of data augmentation stabilizes, the agent proceeds with parameter optimization. This procedure can be repeated a fixed number of times until it meets the stop criteria.
@@ 3. Experiments
In this section, we first introduce the experimental settings, including the datasets, task models, and evaluation metrics. Next, we present the experimental results, covering the performance across multiple trajectory tasks as well as the effectiveness of agent execution. Following that, we perform an ablation study to assess the impact of different components within the proposed framework. Subsequently, we investigate the effect of key parameters for TrajAgent execution. Finally, we conclude with a case study to further illustrate the practical application of our approach.
@@ 3.1. Settings
3.1.1. Dataset
Table 1. Statistics of the Check-in trajectory datasets.
Datasets Foursquare (FSQ) Brightkite(BGK)
Num. Users 463 272
Num. POIs 19870 50061
Num. Trajectories 10632 22208
Table 2. Statistics of the GPS trajectory datasets.
Datasets Porto Chengdu
Sampling Rate 15s 3s
Num. Trajectories 1,710,670 5,819,383
Avg. Trajectory Length (m) 3522.64 2857.81
Avg. Travel Time (s) 724.20 436.12
Latitude Range [41.1401, 41.1859] [30.6529, 30.7277]
Longitude Range [-8.6902, -8.5491] [104.042, 104.129]
3.1.2. Task models
In this work, We adopt various AI models and incorporate them into TrajAgent for solving trajectory-related tasks. According to the type of tasks they deal with, these models can be framed in the following categories:
  • Models for Trajectory Prediction. We consider the classic method RNN (Liu et al., 2016) and well-performing approaches FPMC (Rendle et al., 2010) and DeepMove (Feng et al., 2018) for the task of next location prediction.
  • Models for Trajectory User Linkage. We focus on methods for solving the task of Trajectory User Linkage under check-in trajectories, such as MainTUL (Chen et al., 2022) and DPLink (Feng et al., 2019).
  • Models for Travel Time Estimation. We consider the mainstream model DeepTTE (Wang et al., 2018) to estimate the travel time for a given GPS trajectory.
  • Models for Trajectory Anomaly Detection. We considered the well-performing method GMVSAE (Liu et al., 2020) for detecting anomalous trajectory.
3.1.3. Metrics
In our experiments, Acc@5 is used to measure the accuracy of trajectory prediction and agent execution, Hit@5 is adopted for evaluating the performance of trajectory user linkage task; MAE is employed to compute the error of travel time estimation, while the metric AUC is used to assess the performance of trajectory anomaly detection.
3.2. Main Results
In this section, we evaluate the effectiveness of TrajAgent from two dimensions: 1) Performance of trajectory modelling tasks accomplished through TrajAgent, and 2) Execution efficiency of each stages in TrajAgent.
3.2.1. Performance on Trajectory-related Tasks
Here, we focus on the performances of proposed TrajAgent framework across four types of trajectory tasks: Trajectory Prediction (TP), Trajectory User Linkage (TUL), Travel Time Estimation (TTE) and Trajectory Anomaly Detection (TAD). The first two tasks, TP and TUL, are based on check-in trajectory data, while TTE and TAD utilize GPS trajectory data. The results for these tasks are summarized in Table 3 and Table 4, respectively.
For each task, we report the model performance under four different configurations: the original model (Origin), with Data Augmentation (+DA), with Parameter Optimization (+PO), and with joint optimization (+JO). This setup allows us to assess the individual and combined effects of data augmentation and parameter optimization on the task performance.
Table 3 presents the performance comparison on check-in trajectory tasks TP and TUL. In the TP task on FSQ dataset, DeepMove improves from 0.3422 (Origin) to 0.4018 (+JO), a 17.42% increase. RNN, as a simpler model, experiences the most significant gains, with accuracy increasing by 51.36% through joint optimization. On the BGK dataset, DeepMove continues to perform well, achieving 0.6100 (+JO), marking a 9.52% improvement. Across both datasets, TrajAgent’s optimization techniques consistently enhance model performance, with the FSQ dataset showing particularly notable improvements. For the TUL task, MainTUL and DPLink exhibit consistent improvements under TrajAgent’s optimization strategies. the performance of MainTUL increases from 0.4871 to 0.6121 (+JO) in FSQ, while DPLink achieves a more moderate increase of 6.08%. These results demonstrate the effectiveness of the TrajAgent framework, particularly in automating check-in trajectory tasks. The results also highlight the utility of TrajAgent in terms of integrating data enhancement and parameter optimization strategies for better problem solving.
@@ 3.2.1. Performance on Trajectory-related Tasks
Table 3. Comparison of task performance on check-in trajectories for TrajAgent with different configurations. ‘DA’ represents Data Augmentation, ‘PO’ denotes Parameter Optimization, ‘JO’ indicates Joint Optimization. $\delta$ represents improvements.
Task TP TUL
Model RNN FPMC DeepMove MainTUL DPLink
@@
FSQ Origin 0.1795 0.2939 0.3422 0.4871 0.7551
+DA 0.2667 0.2939 0.4018 0.5973 0.7551
+PO 0.1795 0.3066 0.3422 0.5691 0.7686
+JO 0.2717 0.3066 0.4018 0.6121 0.8010
$\delta$ 51.36% 4.32% 17.42% 25.66% 6.08%
BGK Origin 0.4422 0.5241 0.5570 0.5908 0.8993
+DA 0.5416 0.5275 0.5647 0.6836 0.9613
+PO 0.5022 0.5250 0.6041 0.6683 0.9552
+JO 0.5470 0.5275 0.6100 0.7145 0.9622
$\delta$ 23.70% 0.65% 9.52% 20.94% 6.99%
Task
TTE
TAD
Model DeepTTE GMVSAE
Metrics MAE AUC
Porto origin 8.48 0.9892
+JO 5.85 0.9899
$\delta$ 31.01% minor
Chendu Origin 7.23 0.978
+JO 5.95 0.984
$\delta$ 17.70% 0.61%
3.2.2. Efficiency of Agentic Workflow
We select the trajectory prediction task as an example to demonstrate the efficiency of agentic workflow.
Table 5. Execution efficiency of TrajAgent implemented by different LLMs.
LLM Extraction Processing Data/Model Selection Data Augmentation. Parameter Optim. Joint Optim.
@@ 3.2.2. Efficiency of Agentic Workflow
Table 5 compares the execution efficiency of TrajAgent implemented by different LLMs, across each stages in the trajectory modelling workflow. We can observe that Qwen2-72B and GPT-4o-mini demonstrate the highest success rates (i.e., Succ.) across key stages such as Data Extraction, Processing, and Data/Model Selection, with over 90% success in each. They also excel in Joint Optimization, with success rates of 94% and 96%, resulting in superior Acc@5 scores of 0.4333 and 0.3724 respectively. In contrast, models like Gemma-2-9B and LLama3-8B struggle with lower processing and data selection success rates, which results in reduced overall performance. Their weaker performance in key optimization stages, especially in parameter selection, reflects their limitations in effectively supporting TrajAgent.
These results show that TrajAgent’s efficiency is strongly influenced by the base LLM, with high-performing models like Qwen-72B and GPT-4o-mini significantly enhancing its capabilities. To fully realize TrajAgent’s potential, it is essential to improve LLM performance in data augmentation and parameter optimization, as many LLMs currently struggle in these aspects.
3.3. Ablation Study
@@ 3.4. Parameter Study
In this section, we analyze the effects of two important parameters in the TrajAgent framework: 1) thought step: the number of steps that agent thought before taking action, and 2) memory size: the size of memory units in TrajAgent. To be specific, we vary the values of these parameters to investigate their impact on agent performance, and the results are illustrated in Figure LABEL:fig:param-test(a) and (b). As shown in Figure LABEL:fig:param-test(a), increasing thought steps significantly enhances Joint Optimization (JO) and Data Augmentation (DA). This indicates that giving TrajAgent more reasoning time improves its overall performance. However, the peak performance for Parameter Optimization (PO) and JO at around 10 steps suggests that once sufficient contextual understanding is achieved, further increasing the thought steps provides diminishing returns and may not be beneficial. From Figure LABEL:fig:param-test(b), we find that increasing memory size initially improves performance for JO and DA, with peak accuracy reached at a memory size of 10. This suggests that having sufficient memory allows TrajAgent to store and utilize relevant information effectively. However, beyond this point, performance starts to decline, indicating that excessive memory may introduce unnecessary complexity or irrelevant data, leading to diminished returns for JO and DA. For Parameter Optimization (PO), memory size has little effect, reflecting that this task does not rely as heavily on memory records.
3.5. Case Study

### v2

@@ 2.4.1. Agent Learning via Reasoning
Following Reflexion (Shinn et al., 2024), we design expert optim agent for learning from the experimental records via reasoning. The standard optim agent utilizes the history operation and related results as the feedback to update its action in the next step. Specifically, it works as two stages, including ”think” and ”action”. To support the ”think then action”, it builds a long-term memory for recording all experimental data and a short-term memory for historical actions. In the ”think” stage, optim agent analyzes the long-term memory and meta information of experiment and generating the guidance of action in the short-term memory. In the ”action” stage, optim agent analyze the results in short-term memory to generate the action. Different optimization mechanism utilize the same optim agent with different action space and optimization tips.
Besides, as shown in Figure 1, our framework can also utilize LLM-based agent as the specialized model to complete the trajectory modelling tasks. Thus, we design prompt optimization agent to optimize agent based specialized models. We keep all experimental records in each experiment, including the raw input trajectories, the LLM’s inference results, and the reasoning process. We select the two best-performing trajectories along with their corresponding inference results and reasoning process as few-shot examples, which are then added to the original prompt for a re-run of the experiment. This process of conducting experiments and selecting results is referred to as one iteration. In each iteration, the few-shot examples are updated based on the experiment results.
2.4.2. Model Learning via Training
By collaborating with the high-level agent, the low-level specialized models learn specific trajectory patterns tailored to the target trajectory data and tasks. To enhance performance, we introduce several optimization techniques, including data augmentation, parameter tuning, and joint optimization.
@@ 2.4.2. Model Learning via Training
Data Augmentation. Based on the high-level optim agent, we introduce the specific action space for low-level trajectory data augmentation. Following practice from existing works (Zhou et al., 2024; Dang et al., 2023; Yin et al., 2024), we define a fixed set with ten operators for trajectory data augmentation, e.g., insert, replace, split and so on. During the optimization, the operator set with simple description is provided to the optim agent, it can select any combination of operators with their simple parameters as the action. Then the specialized models are trained with the augmented trajectory data and report performance metrics. The optim agent obtain the feedback information, e.g., performance metrics, from UniEnv to continue update its memory and action.
@@
Parameter Optimization. The action space of parameter optimization is defined based on the parameters of model itself. We define a parameter configuration file for each model, the optim agent reads the configuration file and generates code as the action to update the parameters in it. To better understand the meaning of each parameter, we add comments for each parameter in the file. While the action space is large, this kind of action space is flexible to adapt with any models.
@@
Joint Optimization. Furthermore, we introduce the joint optimization mechanisms to further improve the performance. Due to the different working paradigms, the direct combination of two kinds of optimizations is unsuccessful. We designate the optimization order to prioritize data augmentation first, followed by parameter optimization. This means that once the performance of data augmentation stabilizes, the agent proceeds with parameter optimization. This procedure can be repeated a fixed number of times until it meets the stop criteria.
@@ 3. Experiments
Table 1. Performance of representative methods across five fundamental trajectory modeling tasks. For the nine subtasks, only one model is presented for each. ‘DA’ represents data augmentation, ‘PO’ denotes parameter optimization, ‘PRO’ denotes prompt optimization, ‘JO’ indicates joint optimization, $\delta$ represents performance improvements.
Task Trajectory Prediction Trajectory Recovery Trajecory Classification Trajectory Trajectory
SubTask Next Loc Pre TTE Recovery Map-matching User Linking Intent Prediction Anomaly Generation Representation
Metric Acc@5 MAE MAE $Acc_{m}$ Acc@5 $Acc_{i}$ AUC MAE Acc@5
Dataset FSQ Porto Porto Tencent FSQ Beijing Porto Earthquake FSQ
Models GETNext LLM-ZS MulT-TTE TrajBERT GraphMM S2TUL LIMP GM-VSAE DSTPP CACSR
Origin 0.3720 0.3110 163.12 42.71 0.2014 0.5755 0.745 0.9892 0.4611 0.31
+DA 0.3894 0.6846 0.3369
+PO 0.3995 128.57 27.78 0.2427 0.757 0.9899 0.3584 0.3466
+PRO 0.3225 0.7627
+JO 0.4002 0.3225 128.57 27.78 0.2427 0.7802 0.7627 0.9899 0.3584 0.3472
$\delta$ 7.58% 3.69% 21.18% 34.96% 20.71% 35.57% 2.38% 0 22.27% 12%
@@ 3.1. Settings
Here, we introduce the experimental settings, including the datasets, task models, and evaluation metrics. More detailed information can refer to appendix A.
Data. We utilize the widely used Foursquare (FSQ) (Yang et al., 2016) and Brightkite (BGK) (Cho et al., 2011) in our framework as the default check-in trajectory data. Porto (Kaggle, 2015) and Chengdu (Chuxing, 2018) are integrated in the framework as the default GPS trajectory data. Besides, we use Tencent (Liu et al., 2023) as the road network based methods to support road network based tasks and Beijing (Li et al., 2024) with human labeled intention to support the mobility intention prediction task. Finally, to verify the effectiveness of the whole system, we utilize self-instruct method (Wang et al., 2022) with 5 seed queries to generate 300 user queries as the experiment input.
Models. As shown in Figure 3, our framework supports 18 models spanning 5 core trajectory modeling tasks, which are further categorized into 9 subtasks. The next location prediction task includes FPMC (Rendle et al., 2010), RNN, DeepMove (Feng et al., 2018), GETNext (Yang et al., 2022), LLM-ZS (Beneduce et al., 2024). The travel time estimation task comprises DeepTTE (Wang et al., 2018) and MulT-TTE (Liu et al., 2020a). The trajectory recovery task features TrajBERT (Si et al., 2023), while the map-matching task incorporates DeepMM (Feng et al., 2020) and GraphMM (Liu et al., 2023). The trajectory user linking task includes DPLink (Feng et al., 2019), MainTUL (Chen et al., 2022), and S2TUL (Deng et al., 2023). The mobility intention prediction task is supported by LIMP (Li et al., 2024). The trajectory anomaly detection task employs GM-VSAE (Liu et al., 2020b), and the trajectory generation task includes ActSTD (Yuan et al., 2022) and DSTPP (Yuan et al., 2023). Finally, the trajectory representation method is implemented using CASCR (Gong et al., 2023).
Metrics. We adopt standard practices for each task to select appropriate metrics for evaluating the performance of our framework. The widely used Acc@k metric is employed for next location prediction, map-matching, trajectory user linking, mobility intention prediction, and trajectory representation tasks. The MAE metric is utilized for travel time estimation, trajectory recovery, and trajectory generation tasks. Lastly, the AUC metric is specifically defined for the trajectory anomaly detection task.
3.2. Overall Performance and Generalization Capability of TrajAgent
In this section, we assess the overall performance of TrajAgent across various fundamental trajectory modeling tasks, as summarized in Table 1. Furthermore, to demonstrate the generalization capability of TrajAgent across different datasets and models, we present detailed results on diverse trajectory data using several representative models in Table 2.
As shown in Table 1, we select at least one representative models for 9 trajectory modelling tasks to present the effectiveness of proposed framework. As Table 1 shows, TrajAgent supports a variety of widely-known trajectory models and demonstrates superior performance across multiple trajectory modeling tasks and trajectory datasets. The output of TrajAgent consistently outperforms the original methods, achieving performance gains ranging from 2.28% to 35.57%. For instance, in the next-location prediction task, TrajAgent harnesses its agentic workflow and collaborative learning schema for automatic modeling and optimization, leading to significant performance improvements of various models. Specifically, it enhances the deep learning-based model GETNext by 7.58% and the LLM-based model LLM-ZS by 3.69%. Performance gain in other tasks and models are much larger, for example improvement from 20.71% to 34.96% for trajectory recovery tasks. We observe that the improvement in the trajectory anomaly detection task is minimal, primarily due to the strong performance of the original models on the dataset.
As shown in Table 2, we presents a performance comparison for check-in trajectory tasks, including the next-location prediction task and the trajectory user linking task, across three models and two datasets. The first key observation is the consistent improvement observed across tasks, models, and datasets, which highlights the potential generalization of the proposed framework. For different trajectory tasks, datasets, and models, TrajAgent consistently provides transferable performance improvements, ranging from 6.08% to 35.57%. Additionally, we observe that the performance gap between different models (e.g., the top two models for each task) on specific tasks and datasets significantly narrows from 24.88% to 9.5% following the automatic optimization of TrajAgent, emphasizing the critical role of effective data augmentation and parameter optimization. In other words, with the support of TrajAgent, we can easily obtain more robust and reliable task performance with less relying on the delicate model designs, data processing and parameter selection.
@@ 3.2. Overall Performance and Generalization Capability of TrajAgent
Table 2. Comparison of task performance on check-in trajectories for TrajAgent with different configurations, which demonstrate the generalization of TrajAgent across different tasks, models and datasets. ‘DA’ represents Data Augmentation, ‘PO’ denotes Parameter Optimization, ‘JO’ indicates Joint Optimization. $\delta$ represents improvements.
Task Next Location Prediction Trajectory User Linking
Model RNN DeepMove GETNext MainTUL DPLink S2TUL
@@
FSQ Origin 0.1795 0.3422 0.3720 0.4871 0.7551 0.5755
+DA 0.2667 0.4018 0.3894 0.5973 0.7551 0.6846
+PO 0.1795 0.3422 0.3995 0.5691 0.7686 0.7570
+JO 0.2717 0.4018 0.4002 0.6121 0.8010 0.7802
$\delta$ 51.36% 17.42% 7.58% 25.66% 6.08% 35.57%
BGK Origin 0.4422 0.5570 0.5324 0.5908 0.8993 0.5802
+DA 0.5416 0.5647 0.6026 0.6836 0.9613 0.6903
+PO 0.5022 0.6041 0.6116 0.6683 0.9552 0.7137
+JO 0.5470 0.6100 0.6227 0.7145 0.9622 0.7240
$\delta$ 23.70% 9.52% 16.96% 20.94% 6.99% 24.78%
Table 3. Execution success rates and task performance at each stage of the agentic workflow of TrajAgent across different LLMs. The Acc@5 is obtained by evaluating the next-location prediction task, with DeepMove as the default specialized model.
LLM Extraction Processing Data/Model Selection Data Augmentation Parameter Optim. Joint Optim.
@@
3.3. Ablation Study and Parameter Analysis of the Agentic Workflow
Here, we select the next location prediction task as an example to demonstrate the efficiency of designs of agentic workflow and the effects of various LLMs in Table 3 and Table 4. During the experiment,we select DeepMove as the default specialized model for next location prediction task.
Table 3 compares the execution efficiency of TrajAgent implemented by different LLMs, across each stages in the trajectory modelling workflow. We can observe that Qwen2-72B and GPT-4o-mini demonstrate the highest success rates (i.e., Succ.) across key stages such as Data Extraction, Processing, and Data/Model Selection, with over 90% success in each. They also excel in Joint Optimization, with success rates of 94% and 96%, resulting in superior Acc@5 scores of 0.4333 and 0.3724 respectively. In contrast, models like Gemma-2-9B and LLama3-8B struggle with lower processing and data selection success rates, which results in reduced overall performance. Their weaker performance in key optimization stages, especially in parameter selection, reflects their limitations in effectively supporting TrajAgent. These results show that TrajAgent’s efficiency is strongly influenced by the base LLM, with high-performing models like Qwen-72B and GPT-4o-mini significantly enhancing its capabilities. To fully realize TrajAgent’s potential, it is essential to improve the LLM performance in data augmentation and parameter optimization, as many LLMs currently struggle in these aspects, which is the important future direction of TrajAgent.
@@ 3.3. Ablation Study and Parameter Analysis of the Agentic Workflow
Due to the importance of reflection and memory in the TrajAgent, we analyze the effects of two important related parameters: 1) thought step: the number of steps that agent thought before taking action in reflection, and 2) memory size: the size of memory units in TrajAgent. As shown in Figure LABEL:fig:param-test:thought, increasing thought steps significantly enhances Joint Optimization (JO) and Data Augmentation (DA). This indicates that giving TrajAgent more reasoning time improves its overall performance. However, the peak performance for Parameter Optimization (PO) and JO at around 10 steps suggests that once sufficient contextual understanding is achieved, further increasing the thought steps provides diminishing returns and may not be beneficial. From Figure LABEL:fig:param-test:memory, we find that increasing memory size initially improves performance for JO and DA, with peak accuracy reached at a memory size of 10. This suggests that having sufficient memory allows TrajAgent to store and utilize relevant information effectively. However, beyond this point, performance starts to decline, indicating that excessive memory may introduce unnecessary complexity or irrelevant data, leading to diminished returns for JO and DA. For Parameter Optimization (PO), memory size has little effect, reflecting that this task does not rely as heavily on memory records.
3.4. Comparison with Automated Methods
Figure 5. Compared to Optuna and UrbanLLM, TrajAgent achieves better performance in trajectory user linking tasks with S2TUL as the specialized model. Here, POx represents parameter optimization performed over x steps, while DAx refers to data augmentation conducted over x steps.
Here, we compare TrajAgent with the deep learning-based AutoML method Optuna (Akiba et al., 2019) and the LLM-based automated urban task-solving framework UrbanLLM (Jiang et al., 2024). Using the trajectory user linking task with the S2TUL model as an example, the results are presented in Figure 5. UrbanLLM (represented by the blue line) is designed to directly automate the use of existing models without further optimization, resulting in the lowest performance in Figure 5. Meanwhile, compared to the widely used deep learning-based AutoML method Optuna (represented by the red line), which focuses on parameter optimization, TrajAgent (depicted by the yellow line) achieves superior results with fewer trial-and-error iterations. Furthermore, its performance can be further enhanced through joint optimization combined with automated data augmentation, resulting in a significant performance improvement of over 11.1%.
4. Related Work
4.1. Trajectory Modelling and Analytics
In recent year, trajectory modelling (Zheng, 2015; Chen et al., 2024; Graser et al., 2024) makes great progress on its core research questions, including prediction (Liu et al., 2016; Feng et al., 2018; Shen et al., 2020), classification (Liang et al., 2022; Song et al., 2018), recovery (Xia et al., 2021) and generation (Yuan et al., 2022). Trajectory prediction is to predict the next location or the trip destination of the trajectory. ST-RNN (Liu et al., 2016) is the first work for applying recurrent neural network into the trajectory prediction task. Following this work, lots of works have been done to solve the critical issues of trajectory prediction with deep learning, for example, DeepMove (Feng et al., 2018) for multi-granularity regularity modelling, Flash-back (Yang et al., 2020) for sparse trajectory prediction, GETNext (Yang et al., 2022) for utilizing aggregated mobility pattern. As for the classification problem in trajectory modelling, the most well-known tasks are trajectory mode detection (Jiang et al., 2017) and trajectory user linkage (Gao et al., 2017) task. Various techniques, such as siamese network (Feng et al., 2019), distillation learning (Chen et al., 2022) and transformers (Liang et al., 2022) are successfully applied to discover the pattern among trajectories to solve the classification tasks. Finally, due to the natural limitation of trajectory data quality, trajectory data recovery (Xi et al., 2019; Xia et al., 2021) and generation (Ouyang et al., 2018) are also important tasks in trajectory modelling which are crucial for the practical application in the real world. Thus, follow the the trend of generative modelling, diverse generative models including GAN (Ouyang et al., 2018), diffusion models (Yuan et al., 2022; Wei et al., 2024) are utilized to improve the modelling of human dynamics. While these specific methods accelerate the development of trajectory modelling from different aspects, they can only handle one type of task. In other words, the automated and unified model for all the trajectory modelling task is still missing due to the heterogeneity of tasks and trajectory data. In this paper, we propose to utilize the power of LLM and agent to build a unified model framework for diverse trajectory modelling tasks, which can automatically handle various trajectory data and trajectory modelling tasks without human intervention.
4.2. Large Language Model and Agent
With the advent of ChatGPT (OpenAI, 2024), large language models(LLMs) with extensive commonsense and outstanding reasoning abilities have been widely explored in many domains, such as mathematics (Yao et al., 2022b), question answering (Zhuang et al., 2023), and human-machine interaction (Touvron et al., 2023). Furthermore, agentic framework (Wang et al., 2024; Sumers et al., 2023; Xi et al., 2023) are proposed to enhance the robustness and task solving abilities of LLMs for real-world complex tasks, such as web-navigation (Yao et al., 2022a; Nakano et al., 2021; Zhou et al., 2023) and software development (Hong et al., 2023; Qian et al., 2024; Yang et al., 2024). Following this promising direction, researchers also explore the potential of applying LLM based agent for automatic research experiments (Bogin et al., 2024; Romera-Paredes et al., 2024; Boiko et al., 2023) especially machine learning experiments (Huang et al., 2024; Shen et al., 2024; Zhang et al., 2024; Wu et al., 2024b). For example, MLAgentBench (Huang et al., 2024) designs MLagent to automatically solve simple kaggle competition with ReAct mechanism in terms of success rate. DSBench (Zhang et al., 2024) verifies the LLM based agent for solving classic data science tasks with tabular data as input. However, trajectory data are more complicated than the tabular data or text data and different trajectory modelling tasks also exhibit diverse problem settings and models, which are too challenging for these existing agents designed for general machine learning tasks with simple data input and interface. Recently, UrbanLLM (Jiang et al., 2024) follows the idea of HuggingGPT (Shen et al., 2024) to manage the existing specialized spatial-temporal models and APIs for tackling urban related tasks, which covers some trajectory modelling tasks. However, existing specialized spatial-temporal AI models usually designed for specific task with specific data and format, directly utilizing these models can suffer significant performance drop even can not generate any make sense results. In this paper, different from general machine learning agent and UrbanLLM for static spatial-temporal model, our proposed agentic framework is designed for unified trajectory modelling which can provide automatically model training and optimization for various trajectory data and tasks.
5. Conclusion
In this paper, we propose TrajAgent, an LLM-based agentic framework for automated trajectory modeling. Supported by UniEnv, which provides a unified data and model interface, and collaborative learning schema for joint performance optimization, TrajAgent can automatically identify and train the appropriate model, delivering competitive performance across a range of trajectory modeling tasks. TrajAgent introduces a new paradigm for unified trajectory modeling across diverse datasets and tasks, paving the way for advancements in trajectory modeling applications.
Appendix A Appendix
A.1. Datasets
  • Tencent (Liu et al., 2023): This dataset includes both a road network and a set of vehicle trajectories collected in northeastern Beijing. The road network consists of 8.5K road segments and 15K edges, and the trajectory dataset contains 64K vehicle trajectories, each sampled at 15-second intervals.
  • Beijing (Li et al., 2024): This dataset contains check-in records collected from a popular social networking platform. It spans a period from late September 2019 to late November 2019.It also contains intent labels annotated by human for a small dataset.
Table 5. Statistics of the Check-in trajectory datasets.
Datasets Foursquare (FSQ) Brightkite(BGK) Beijing
Num. Users 463 272 1566
Num. POIs 19870 50061 5919
Num. Trajectories 10632 22208 744813
Table 6. Statistics of the GPS trajectory datasets.
Datasets Porto Chengdu Tencent
Sampling Rate 15s 3s 15s
Num. Traj. 1,710,670 5,819,383 10,000
Avg. Traj. Length (m) 3522.64 2857.81 2492.01
Avg. Travel Time (s) 724.20 436.12 13903.78
Latitude Range [41.1401, 41.1859] [30.6529, 30.7277] [40.0224, 40.0930]
Longitude Range [-8.6902, -8.5491] [104.042, 104.129] [116.265, 116.349]
A.2. Models
As introduced in Figure 3, we adopt various deep learning based and LLM based models and incorporate them into TrajAgent for solving trajectory-related tasks. According to the type of tasks they deal with, these models can be framed in the following categories:
  • Next Location Prediction: RNN (Liu et al., 2016), attention-based method like DeepMove (Feng et al., 2018), well-performed graph-based method like GETNext (Yang et al., 2022) and two recent proposed LLM-based methods LLM-ZS (Beneduce et al., 2024) and LLMMove (Feng et al., 2024).
  • Travel Time Estimation: DeepTTE (Wang et al., 2018) and MulT-TTE (Liu et al., 2020a) to estimate the travel time for a given GPS trajectory.
  • Trajectory User Linkage: widely-used DPLink (Feng et al., 2019), MainTUL (Chen et al., 2022) and S2TUL (Deng et al., 2023) are considered. We modified DPLink’s training approach by using publicly available sparse trajectory datasets instead of heterogeneous mobility datasets for training.
  • Travel Intent Prediction: LIMP (Li et al., 2024), which leverages the commonsense reasoning capabilities of LLMs for mobility intention inference.
  • Trajectory Anomaly Detection: we consider the well-performing method GMVSAE (Liu et al., 2020b), which represent different types of normal trajectories in a continuous latent space.
  • Trajectory Generation: ActSTD (Yuan et al., 2022), which capture the spatiotemporal dynamics underlying trajectories by leveraging neural differential equations and DSTPP (Yuan et al., 2023), which defines spatial temporal point process for trajectory generation.
  • Trajectory Recovery: TrajBERT (Si et al., 2023), which encode trajectory as sentence and train a BERT model to get representations of trajectories.
  • Trajectory Map Matching: DeepMM (Feng et al., 2020), which proposes a data augmentation approach for map-based trajectory data, and GraphMM (Liu et al., 2023), which leverages a graph-based framework to extract features from map-based trajectory data, are considered.
  • Trajectory Representation: CASCSR (Gong et al., 2023), which use contrastive learning method to learn trajectory representations for downstream tasks.
A.3. Metrics
  • $Acc_{m}$: $Acc_{m}$ is the evaluation metric which computes the average matching degree of all trajectories.For each trajectory, its matching degree is the ratio of the number of matching road segments to the number of all road segments.
  • $Acc_{i}$: $Acc_{i}$ is used to measure the accuracy of trajectory intention inference. It is the ratio of the number of matching predicted intention of each check-in to the total number of check-ins.
  • JSD: Jensen–Shannon divergence (JSD) measures the discrepancy of distributions between the generated data and real-world data. Lower JSD denotes a closer match to the statistical characteristics and thus indicates a better generation result.
In our experiments, Acc@5 is used to measure the accuracy of trajectory prediction and agent execution, it is also used to measure the downstream applications of trajectory representation.$Acc_{m}$ is designed for measuring the performance of map matching task.$Acc_{i}$ is designed for measuring the performance of travel intention prediction task. Hit@5 is adopted for evaluating the performance of trajectory user linkage task; MAE is employed to compute the error of travel time estimation and trajectory recovery, while the metric AUC is used to assess the performance of trajectory anomaly detection.JSD and RMSE are used to measure the prediction error of the spatiotemporal domain in trajectory generation task.
A.4. Additional results on GPS trajectory data
Task TTE Anomaly Recovery
Model DeepTTE MultTTE GMVSAE TrajBERT
Metrics MAE AUC MAE
Porto origin 8.48 129.35 0.9892 13.6667
+JO 5.85 109.85 0.9899 8.0290
$\delta$ 31.01% 15.08% minor 41.25%
Chendu Origin 7.23 166.29 0.978 54.8060
+JO 5.95 128.57 0.984 29.8134
$\delta$ 17.70% 22.68% 0.61% 54.39%
A.5. Case Study